# InvisibleInk: High-Utility and Low-Cost Text Generation with Differential Privacy

### v1

@@ 5.2 Experimental Results
We now compare the privacy-utility-compute tradeoffs of InvisibleInk with the baselines.
Fig. 4 compares InvisibleInk with the baselines at a target level of $(\varepsilon=10,\delta=10^{-6})$-DP. See Tab. 2 and §E for evaluations at other $\varepsilon$ values.
We find that InvisibleInk with Top-$k+$ sampling consistently produces the highest quality text with the longest generations (and the most medical named entities for MIMIC) at smaller compute budgets across datasets. Consider the results on MIMIC, for example: InvisibleInk can achieve a MAUVE score of $\approx 71\text{-}74\%$ with 8-16 inference calls per token, while Amin et al. [12] achieve $\approx 70\text{-}72\%$ at with 64-128 inference calls: this is an 8$\times$ improvement in computation cost.
InvisibleInk also produces consistently more medical named entities than Amin et al. [12] despite producing generations of nearly the same length. Thus, InvisibleInk is able to better capture the domain-specific information in the sensitive references and produce qualitatively better text.
AdaPMixED [18] Amin et al. [12] InvisibleInk$\;\;{(k=|V|)}$ InvisibleInk$\;\;{(k=50)}$ $\bm{B}$ $\bm{7}$ $\bm{31}$ $\bm{127}$ $\bm{31}$ $\bm{63}$ $\bm{127}$ $\bm{3}$ $\bm{7}$ $\bm{15}$ $\bm{3}$ $\bm{7}$ $\bm{15}$ $\varepsilon=1$ $58.7$ $58.7$ $58.5$ INF INF INF $62.9$ $63.6$ $64.4$ 67.867.8{67.8}67.8 $\mathbf{68.8}$ 65.965.9{65.9}65.9 $\varepsilon=3$ $58.3$ $58.8$ $61.5$ INF INF $55.4$ $62.1$ $63.3$ $64.5$ 66.866.8{66.8}66.8 68.268.2{68.2}68.2 $\mathbf{71.0}$ $\varepsilon=5$ $60.2$ $58.7$ TLE INF INF 70.070.0{70.0}70.0 $64.4$ $63.5$ $67.9$ 66.666.6{66.6}66.6 $68.3$ $\mathbf{70.5}$ $\varepsilon=10$ $60.2$ $59.2$ TLE INF 71.971.9{71.9}71.9 $69.8$ $65.0$ $63.9$ $66.8$ $70.2$ $71.0$ $\mathbf{74.2}$
@@
Tab. 2 shows that InvisibleInk exhibits graceful degradation in utility when operating under extremely tight privacy budgets or compute constraints, where baselines fail. Specifically, Amin et al. [12]’s method uses up all its privacy budget for the SVT term and cannot produce any text at (a) $B\leq 31$, (b) $\varepsilon\leq 5$ and $B=63$, and (c) $\varepsilon=1$ and $B=127$; see §E. Meanwhile, AdaPMixED extinguishes its data-dependent privacy budget after generating as few as 10 tokens. Both baselines thus require large batch sizes to operate meaningfully, rendering them impractical in compute-constrained environments.
Comparison with API-Access Methods. Unlike InvisibleInk and other baselines considered in this paper, Augmented Private Evolution (AugPE) [36] assumes only API access to the LLMs. This more restrictive setting naturally makes privatization harder. AugPE uses an evolutionary algorithm: it iteratively uses LLM-based paraphrasing to construct candidate generations, and maintains the “more appropriate” generations at each step by constructing privatized (using Gaussian mechanism) nearest neighbor histograms (cf. §B.3.3) of the candidate synthetic generations at every iteration.
In particular, AugPE makes two types of LLM inference calls: generation and paraphrasing. While paraphrasing can be implemented with calls to the same decoder-only LLMs that are also used for generation, it is also possible to use cheaper paraphrasing models or masked language models. Thus, we cannot directly compare the computational cost of InvisibleInk to AugPE in an apples-to-apples comparison based on the number of next-token inference calls. Instead, we directly use the wall-clock runtime of both methods as a measure of computation cost. Similar to other methods, we also allow AugPE an advantage of $\sqrt{2}$ in the sensitivity in the replace-by-null adjacency (see §B.3.3 for details). The results are given in Fig. 5: additional results and detailed generation settings are discussed in §E.1.
Figure 5: Utility vs Computational Cost plots for InvisibleInk and AugPE for $\varepsilon=10$ for 1000 synthetic texts generated for the MIMIC dataset. We compare utility using a variety of metrics — InvisibleInk outperforms AugPE across all settings and evaluation metrics. Wall-clock run time is used as a proxy for computational cost. We report results for $B+1=4,8,16,32$ next-token inference calls per generated token and $T_{\mathsf{AugPE}}=1,3,5,10$ for InvisibleInk and AugPE, respectively.
@@
We find that InvisibleInk completely outperforms AugPE across privacy budget settings and evaluation metrics in Fig. 5. The overall quality of synthetic text generated by AugPE, even for the most relaxed settings (in terms of both compute and privacy budgets), falls short of that of the most compute and privacy-constrained generations of InvisibleInk. We found in our preliminary experiments that AugPE showed competitive (with InvisibleInk) performance on the Yelp dataset. This can likely be attributed to the LLM-based paraphrasing step working best when the dataset of private references is well-represented in the pre-training dataset. We remark that the dominance of InvisibleInk over AugPE is not surprising, as the API access setting of AugPE is much more restrictive compared to the whitebox setting of InvisibleInk.
5.3 Understanding InvisibleInk: Ablations and Scaling
We examine the importance of each of the two key components—DClip and the Top-$k+$ sampling—of InvisibleInk. Fig. 3 shows that the smaller active range of ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ allows us to use a smaller clip norm to keep the same level of distortions. This directly allows us to use smaller batch sizes $B$ to keep $\tau\propto C/B$ around 1. We see from Fig. 4 that this translates to an approx. 2×2$\times$2 × improvement in computation cost at the same privacy-utility levels.
@@ 5.3 Understanding InvisibleInk: Ablations and Scaling
Figure 6: Truncated decoding offers significant benefits: MAUVE scores for DP synthetic text generation of $1000$ samples from the MIMIC Dataset for various privacy budgets using TinyLLaMA-1B and LLaMA3.2-1B models for various top-$k$ thresholds at a temperature $\tau=1.2$ and batch size $B=7$.
Fig. 6 shows us that Top-$k+$ decoding offers improvements over using the full vocabulary ($0.71$ vs. $0.64$ at $\varepsilon=10$ for TinyLlama). This distinction is more pronounced for Llama-3.2 1B model ($0.81$ vs. $0.59$ MAUVE at $\varepsilon=10$), which is of similar size but has a 4$\times$ larger vocabulary ($128K$ vs. $32K$).
We also find that Top-$k+$ expansion leads to $7.8\pm 3.4$ tokens being introduced by the expansion when $k=10$ and $\varepsilon=10$ using a batch size $B=7$. This expanded vocabulary accounts for $1.75\pm 0.68\%$ of the generated tokens. See §E for further details. In particular, §E.3 gives additional ablations on the effect of the temperature parameter and clip norm parameters.
Scaling InvisibleInk to Larger Models. The utility scores (MAUVE) for synthetic text generation of $n=500$ generations using a larger LLaMA3 8B model is presented in Tab. 3.888 These results are not directly comparable to the previous results of the 1B models as they used $n=1000$ generations (compared to $n=500$ here), and thus different MAUVE hyperparameters.
The results are qualitatively similar to the ones observed for smaller models, where $k=50$ consistently shows better performance compared to other Top-$k+$ parameters. This highlights the scalability of InvisibleInk to larger models, and by extension, its suitability for high-quality text generation.
$\bm{\varepsilon}$ $\bm{k{=}10}$ $\bm{k{=}50}$ $\bm{k{=}100}$
1.0 59.79 62.76 61.50
3.0 60.55 64.73 62.26
5.0 61.03 65.78 62.41
10.0 59.86 64.93 64.70
[…]
@@ 5.4 Practical Recommendations
We recommend tuning a temperature parameter $\tau\approx 1$ (generally in the range of $0.9$ to $1.2$), if permitted by the computation budget. A reasonable heuristic to avoid tuning is to directly set $\tau=1$, as it achieves competitive performance. Then, we set a clip norm $C=B\tau/\sqrt{2\rho_{\mathsf{seq}}/T}$, where the batch size $B$ fixes the computation budget, $\rho_{\mathsf{seq}}$-zCDP is the target privacy guarantee, and $T$ is the maximum token budget. Finally, we must tune a free parameter $k$ for Top-$k+$ sampling. We recommend tuning $k$ following best practices for non-private decoding (e.g., $k\in[50,100]$ for open-ended settings and $k$ smaller for more directed settings such as reasoning).
@@ 6 Conclusion
In this work, we introduce InvisibleInk, a scalable framework for long-form text generation that provides rigorous differential privacy guarantees with respect to sensitive references. As LLM-based generation increasingly powers paradigms such as retrieval-augmented generation (RAG) and inference-time scaling, ensuring privacy in the presence of sensitive inputs remains a critical challenge.
InvisibleInk casts next-token sampling as an instance of the exponential mechanism over LLM logits and introduces two key innovations. First, it reduces privacy cost by isolating and clipping only the sensitive portion of the logits, measured relative to public logits. Second, it improves utility by sampling from a tight superset of the top-$k$ private tokens. Empirical results demonstrate an 8×8$\times$8 × reduction in compute cost compared to state-of-the-art baselines, while preserving utility across privacy levels. Notably, InvisibleInk enables private long-form generation at less than 8$\times$ the cost of non-private decoding, offering a practical and efficient solution for privacy-preserving language generation.
@@
Second, InvisibleInk provably protects against privacy leakage through inference, but not in the training or fine-tuning stage. Leakage could still be possible through data contamination or backdoor adversarial attacks. Furthermore, even in the case of inference, we tackle example-level DP only. However, user-level privacy leakage is also possible [e.g. 78], and extending InvisibleInk to user-level DP is an interesting direction.
Finally, while InvisibleInk drastically improves the computational cost of private text generation, the 8×8$\times$8 × overhead over non-private generation could still be restrictive with large frontier models under extremely tight privacy budgets. Pushing the envelope of the privacy-utility-compute tradeoffs of text generation is a fruitful avenue for future research.
@@
However, the existence of formal privacy guarantees may lead to overconfidence or misuse, particularly if system integrators overlook practical limitations such as implementation errors, misconfigured parameters, or unrealistic threat models [e.g., 79, 80]. Differential privacy is a powerful but narrow tool—it provides provable anonymization but is only one component in an end-to-end system. The practical use of such a system must be considered in a holistic manner, along with clear communication of failure modes.
Acknowledgements. The authors thank Ambreesh Parthasarathy, Krithi Shailya, Danish Pruthi, and Balaraman Ravindran for fruitful discussions and helpful comments. This research was supported by the Post-Baccalaureate Fellowship at the Centre for Responsible AI (CeRAI, IIT Madras), the startup compute grant of the Wadhwani School of Data Science & AI (IIT Madras), and faculty research awards.
@@ Appendix A Notation
We summarize our main notation in Tab. 4.
@@
${\bm{x}}_{<t}$ $=(x_{1},\ldots,x_{t-1})\in V^{t-1}$, all previously generated tokens before time-step $t$-token
@@
$N$ $=|\mathcal{D}|$, dataset size
@@
$\mathsf{DClip}_{C}({\bm{\phi}},{\bm{\phi}}_{\mathsf{pub}})$ $={\bm{\phi}}_{\mathsf{pub}}+\mathsf{clip}_{C}({\bm{\phi}}-{\bm{\phi}}_{\mathsf{pub}})$, our proposed clipping function that clips the difference with public logits
@@
$B$ $=|{\bm{R}}|$, batch size; controls the computational cost of InvisibleInk

### v2

@@ 5.1 Experimental Results
We compare the privacy-utility-compute tradeoffs of InvisibleInk with the baselines at a target level of $(\varepsilon=10,\delta=10^{-6})$-DP in Fig.˜4. See Tab.˜2 for evaluations at other $\varepsilon$ values (with $\delta=10^{-6}$).
@@
InvisibleInk(k=100)\;{(k=100)} InvisibleInk(k=|V|)\;{(k=|V|)} Amin et al. [13] AdaPMixED [19] $\bm{B}$ $\bm{3}$ $\bm{7}$ $\bm{15}$ $\bm{31}$ $\bm{3}$ $\bm{7}$ $\bm{15}$ $\bm{31}$ $\bm{63}$ $\bm{127}$ $\bm{7}$ $\bm{31}$ $\bm{127}$ $\varepsilon=1$ $68.3_{\;\text{0.1}}$} $66.9_{\;\text{1.1}}$} $67.9_{\;\text{1.5}}$} $\mathbf{69.2_{\;\text{1.3}}}$ $67.3_{\;\text{0.9}}$ $67.6_{\;\text{0.9}}$ $68.2_{\;\text{1.4}}$ INF INF INF $58.8_{\;\text{0.3}}$ $56.8_{\;\text{0.1}}$ $59.2_{\;\text{0.2}}$ $\varepsilon=3$ $67.4_{\;\text{1.1}}$} $69.0_{\;\text{0.7}}$} $69.7_{\;\text{0.6}}$} $\mathbf{70.9_{\;\text{1.2}}}$ $66.9_{\;\text{2.0}}$ $69.1_{\;\text{1.3}}$ $69.1_{\;\text{0.7}}$ INF INF $55.2_{\;\text{0.0}}$ $58.8_{\;\text{0.1}}$ $58.9_{\;\text{0.3}}$ $64.7_{\;\text{1.0}}$ $\varepsilon=5$ $67.7_{\;\text{0.1}}$} $70.3_{\;\text{1.0}}$ $69.7_{\;\text{0.9}}$} $\mathbf{73.5_{\;\text{1.0}}}$ $68.6_{\;\text{0.5}}$ $68.9_{\;\text{0.7}}$ $69.0_{\;\text{0.7}}$ INF INF $69.5_{\;\text{0.0}}$} $59.4_{\;\text{0.3}}$ $59.8_{\;\text{0.4}}$ $69.9_{\;\text{0.9}}$ $\varepsilon=10$ $68.4_{\;\text{1.4}}$ $72.0_{\;\text{0.2}}$ $73.7_{\;\text{0.4}}$} $\mathbf{75.0_{\;\text{1.5}}}$ $68.3_{\;\text{0.9}}$ $70.4_{\;\text{1.3}}$ $72.4_{\;\text{1.4}}$ INF $68.9_{\;\text{0.5}}$} $70.1_{\;\text{1.3}}$ $59.2_{\;\text{0.5}}$ $58.9_{\;\text{0.5}}$ TLE
Table 2: MAUVE ($\%$) scores, reported with mean and $95\%$ confidence intervals (see §D.6) over 33 runs, for $1000$ synthetic generations of the MIMIC dataset using TinyLLaMA. Synthetic generation without any private references ($\varepsilon=0$) yields a MAUVE of $\mathbf{68.56}\%$ (best hyperparameters). InvisibleInk outperforms every private baseline with $\approx\bm{8\times}$ smaller batch size. The results for InvisibleInk with $k=|V|$ are reported for $\tau=1.0$ and for $k=100$ are reported for $\tau=1.1$. AdaPMixED [19] sometimes exceeds a wall clock time of $36$ hours (denoted TLE- Time Limit Exceeded); Amin et al. [13]’s method fails to generate any synthetic text for small privacy/compute budgets (denoted INF- infeasible). The top two scores for every $\varepsilon$ are highlighted.
@@
Tab.˜2 shows that InvisibleInk exhibits graceful degradation in utility when operating under tight privacy budgets or compute constraints, where baselines fail. Specifically, Amin et al. [13]’s method uses up all its privacy budget for the SVT term and cannot produce any text at (a) $B\leq 31$, (b) $\varepsilon\leq 5$ and $B\leq 63$, and (c) $\varepsilon=1$ and $B\leq 127$; see §E. Meanwhile, AdaPMixED extinguishes its data-dependent privacy budget after generating as few as 10 tokens for $B=8$. Both baselines require large batch sizes to operate meaningfully, rendering them impractical in compute-constrained environments. We discuss the full sets of hyperparameters explored, in §D.5.
@@
Comparison with API-Access Methods. We find that InvisibleInk completely outperforms the API-access based method AugPE [41] across privacy budget settings and evaluation metrics in Fig.˜5. We remark that this is expected due to the more restrictive API-access setting of AugPE. Interestingly, we found in preliminary experiments that AugPE showed competitive (with InvisibleInk) performance on the Yelp dataset. This can likely be attributed to the LLM-based paraphrasing step working best when the dataset of private references is well-represented in the pre-training data. See §E.1, for more details and additional results.
Figure 5: InvisibleInk outperforms the API-access method of AugPE [41] across all settings: Utility vs Compute plots (avg. over 33 runs with $95\%$ confidence intervals) for InvisibleInk and AugPE for $\varepsilon=10$ for 1000 synthetic texts generated for the MIMIC dataset. Wall-clock run time is used as a proxy for computational cost. We report results for $B+1=4,8,16,32$ for InvisibleInk and and $T_{\mathsf{AugPE}}=1,3,5,10$ for AugPE.
Downstream Utility of Synthetic Text. We evaluate the synthetic data generated from the Yelp dataset on a downstream classification task. We finetune a RoBERTa model [89] for sequence classification on a dataset of $500$ generated synthetic texts from the Yelp dataset into $50$ classes, obtained as all combinations of $5$ review score labels and $10$ business category labels. We note that since the synthetic data generation is conditioned on these classes, the class label information is also available in the synthetic data.
We evaluate the resulting classifier on the Yelp test set. Tab.˜3 reports the 50-class top-1 and top-5 accuracy (resp. columns “Accuracy” and “Top-5 Acc.”), 10-class accuracy over categories (column “Cat. Acc.”), 5-class accuracy over score labels (column “Score Acc.”), and the $L_{1}$ distance between the predicted and true numeric scores (column “Score $L_{1}$”). We finetune the RoBERTa model [89] for $15$ epochs with an AdamW optimizer [90] with a batch size of $20$, a learning rate of $2\times 10^{-5}$, a weight decay of $0.01$ and $10\%$ warmup steps.
The results presented in Tab.˜3 compare the downstream utility metrics for text generated by various methods with a fixed privacy budget of $\varepsilon=10$. InvisibleInk outperforms all baseline methods by a large margin at a significantly smaller computation cost, even when measured in terms of this downstream task utility.
Method Batch Size $B$ Accuracy ($\uparrow$) Top-5 Acc. ($\uparrow$) Cat. Acc. ($\uparrow$) Score Acc. ($\uparrow$) Score $L_{1}$ ($\downarrow$) InvisibleInk $7$ 32.98 72.16 64.90 52.2 0.652 Amin et al. [13] $127$ 29.44 64.56 60.18 46.64 0.748 AdaPMixED [19] $31$ 7.44 34.72 56.82 15.44 1.858
Table 3: Downstream utility of synthetic text. Multiple evaluation metrics task for a sequence classification task using a RoBERTa [89] model finetuned on synthetic sequences generated from the Yelp dataset with privacy budget $\varepsilon=10$. InvisibleInk outperforms the baselines significantly.
We note that measuring utility in downstream tasks is not a reliable indicator of overall text quality, coherence, or distributional similarity to the original dataset. Several prior works on DP text generation [48, 47, 13, 19] rely heavily on such metrics to evaluate generated text. However, it is possible for poor-quality synthetic text to “hack" these metrics, especially in the case of sequence classification by relying on keywords related to the target category. This can be observed in the case of AdaPMixED, where sequences with an average length of $\leq 30$ tokens (see Fig.˜4) perform significantly better than random: $\approx 7\%$ (vs. $2\%$ at random) for overall classification accuracy, $\approx 56\%$ (vs. $10\%$ at random) for category prediction accuracy, and $\approx 1.8$ (vs. $2.0$ at random) for average $L_{1}$ distance (lower is better) between predicted and true scores. We also note that in the case of open-ended text generation, such notions of downstream utility are not directly applicable. Thus, we adopt a combination of metrics to evaluate DP synthetic text.
5.2 Understanding InvisibleInk: Ablations, Scaling and Additional Experiments
@@
Figure 6: Truncated decoding offers significant benefits: MAUVE scores (avg. over 3 runs with $95\%$ confidence intervals) for DP synthetic text generation of $1000$ samples from the MIMIC Dataset using TinyLLaMA-1B and LLaMA3.2-1B models at a temperature $\tau=1.1$ and batch size $B=7$.
We examine the importance of the two key components—DClip and the Top-$k+$ sampling—of InvisibleInk. Fig.˜3 shows that the smaller active range of ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ allows us to use a smaller $C$, when clipping using DClip, to maintain similar levels of distortion. Thus, we can use smaller batch sizes $B$ to keep $\tau\propto C/B\approx 1$, translating into an $\approx 8\times$ improvement (see Fig.˜4) in computation cost at the same privacy-utility levels. Fig.˜6 shows us that Top-$k+$ decoding offers significant improvements (in MAUVE scores) over using the full vocabulary ($72\%$ vs. $65\%$ at $\varepsilon=10$ for TinyLlama). This distinction is more pronounced for the similar-sized Llama-3.2 1B model ($76\%$ vs. $62\%$ at $\varepsilon=10$), with a $4\times$ larger vocabulary ($128K$ vs. $32K$).
Additional Experiments. In addition to these, we also investigate the following:
  • We present additional statistics of Top-$k+$ sampling from the expansion set ($V_{k}^{+}\setminus V_{k}$), in §E.2 and show empirically that $V_{k}^{+}$ is a tight superset of $V_{k}$.
  • We also highlight the scalability of InvisibleInk to larger models by generating synthetic text using a LLaMA3 8B model at low batch sizes (incompatible with baselines) in §E.3.
  • §E.4 gives additional ablations on the effect of the temperature, clip norm, and top-$k$ parameters. In particular, we show that setting the temperature $\tau\in[1.0,1.1]$ and the top-$k$ parameter $k\approx 100$ yields near-optimal generations across settings.
  • As noted previously, a comparison of the performance of AdaPMixED for replace-by-null and add-or-remove adjacency is also presented in §E.5 to show that the quantitative results for either notion follows similar trends.
@@ 5.3 Practical Recommendations
We recommend tuning a temperature parameter $\tau\approx 1$ (generally in the range of $0.9$ to $1.2$), if permitted by the computation budget. A reasonable heuristic to avoid tuning is to directly set $\tau=1$, as it achieves competitive performance. Then, we set a clip norm $C=B\tau/\sqrt{2\rho_{\mathsf{seq}}/T}$, where the batch size $B$ fixes the computation budget, $\rho_{\mathsf{seq}}$-zCDP is the target privacy guarantee, and $T$ is the maximum token budget. Finally, we must tune a free parameter $k$ for Top-$k+$ sampling. We recommend tuning $k$ following best practices for non-private decoding (e.g., $k\in[50,100]$ for open-ended settings and smaller for more directed settings such as reasoning).
@@ 6 Conclusion
We introduce InvisibleInk, a scalable framework for long-form text generation that provides rigorous differential privacy guarantees with respect to sensitive references.
InvisibleInk casts next-token sampling as an instance of the exponential mechanism over LLM logits and introduces two key innovations. First, it reduces privacy cost by isolating and clipping only the sensitive portion of the logits, measured relative to public logits. Second, it improves utility by sampling from a tight superset of the top-$k$ private tokens. Empirical results demonstrate an $8\times$ (or more) reduction in compute cost compared to state-of-the-art baselines, while preserving utility, across privacy levels. Notably, InvisibleInk enables private long-form generation at less than $4\text{-}8\times$ the cost of non-private decoding, offering a practical and efficient solution for privacy-preserving language generation. While InvisibleInk significantly reduces the computational complexity of generating DP text, it still has some limitations that we now discuss.
@@
Second, InvisibleInk provably protects against privacy leakage through inference, but not in the training or fine-tuning stage. Leakage could still be possible through data contamination or backdoor adversarial attacks. Furthermore, even in the case of inference, we tackle example-level DP only. However, user-level privacy leakage is also possible [e.g. 91], and extending InvisibleInk to user-level DP is an interesting direction.
Finally, while InvisibleInk drastically improves the computational cost of private text generation, the $8\times$ overhead over non-private generation could still be restrictive with large frontier models under extremely tight privacy budgets. Pushing the envelope of the privacy-utility-compute tradeoffs of text generation is a fruitful avenue for future research.
@@
However, the existence of formal privacy guarantees may lead to overconfidence or misuse, particularly if system integrators overlook practical limitations such as implementation errors, misconfigured parameters, or unrealistic threat models [e.g., 82, 83]. Differential privacy is a powerful but narrow tool: it provides provable anonymization but is only one component in an end-to-end system. The practical use of such a system must be considered in a holistic manner, along with clear communication of failure modes.
Acknowledgments
The authors thank Ambreesh Parthasarathy, Krithi Shailya, Danish Pruthi, and Balaraman Ravindran for fruitful discussions. VV is supported by the Post-Baccalaureate Fellowship at the Centre for Responsible AI (CeRAI), IIT Madras. KP is grateful for funding from ANRF’s PM ECRG (ANRF/ECRG/2024/004321/PMS), Schmidt Sciences’ AI2050 program, the startup compute grant of the Wadhwani School of Data Science & AI (WSAI), IIT Madras, and faculty research awards.
@@ Appendix A Notation
We summarize our main notation in Tab.˜4.
@@
${\bm{x}}_{<t}$ $(x_{1},\ldots,x_{t-1})\in V^{t-1}$, all previously generated tokens before time-step $t$-token
@@
NN $|\mathcal{D}|$, dataset size
@@
$\mathsf{DClip}_{C}({\bm{\phi}},{\bm{\phi}}_{\mathsf{pub}})$ ${\bm{\phi}}_{\mathsf{pub}}+\mathsf{clip}_{C}({\bm{\phi}}-{\bm{\phi}}_{\mathsf{pub}})$, our proposed clipping function that clips the difference with public logits
@@
$B$ $|{\bm{R}}|$, batch size; controls the computational cost of InvisibleInk